OpenStudy (anonymous):

Passengers at International Jetport Year Number of Passengers (millions) 1980 33.4 1985 42.1 1989 47.8 1995 58.2 1998 65.7 2005 78.4 Based on the data shown, predict about how many passengers, in millions, used International Jetport in the year 2011? A) 84 B) 87 C) 89 D) 91

OpenStudy (anonymous):

arithmetic?

OpenStudy (anonymous):

algebra

OpenStudy (kropot72):

The first step is to find the average increase in passengers per year. This can be found by taking the number of passengers in 2005 (78.4 million) and subtracting the number of passengers in 1980 (33.4 million). The result is then divided by the time span (2005 - 1980 = 25). \[Average\ increase\ per\ year=\frac{78.4-33.4}{25}\ million\] The predicted number of passengers in 2011 is found by multiplying the average increase by the number of years between 2005 and 2011 (which is 6 years) and adding the result to the number of passengers in 2005 (78.4 million). \[Predicted\ number\ 2011=78.4+\frac{6(78.4-33.4)}{25}=you\ can\ calculate\]

OpenStudy (anonymous):

what would the answer be

OpenStudy (kropot72):

@destinyb4302 Do you have access to a calculator?